OpenStudy (anonymous):

Let f(x) = x^2 - 81. Find f^-1(x). Thanks! :-)

OpenStudy (anonymous):

We can solve for x f(x) = x^2 - 81 f(x) + 81 = x^2 \[\sqrt{ (f(x) + 81)}\](f(x) + 81) = x Your inverse function is f-1(x) = \[\sqrt{(x+81)}\]

OpenStudy (anonymous):

Thank you! That's the answer I originally got! I just wanted to confirm.